OpenStudy (anonymous):

how do u subtract fractions? example 5/21 - 1/6 is it similar to adding? finding the same denominator??

OpenStudy (unklerhaukus):

yep exactly

OpenStudy (unklerhaukus):

5/21 -1/6 30/126 - 21/126 =(30-21)/126

OpenStudy (unklerhaukus):

=9/126

OpenStudy (anonymous):

i used the online calculator and it gave me an answer of 1/14

OpenStudy (unklerhaukus):

9/126=1/14

OpenStudy (anonymous):

how does 14 go into 126? my teacher showed me how to do it a weird way. reduce 126 into 14 x 9 and for top 1 x 9 cross out the same numbers to get 1/14.. if that makes any sence

OpenStudy (unklerhaukus):

makes sense to me, does it make sense to you?

OpenStudy (anonymous):

sometimes cuz i use the same method on alot of probs and some of them are wrong. its confusing

OpenStudy (unklerhaukus):

well if you have a fraction with big numbers in it, try to break the numerator and the denominator into factors, if factors are common they can cancel

OpenStudy (anonymous):

okay, ill keep that in my notes. thanks for the help!